Regarding the HDI, a country in the year 2000 was in the 69th position and in 2001 it was 65th. The new position in relation to the previous one had a percentage increase of approximately:
Answer by Theo(13342)

i think this depends on the number of countries in the index, although i'm not exactly sure how it's calculated.

for example, if there are 100 countries in the index, then the country that is in the 69th position has 31 countries with scores less than it.

when the same country is in the 65th position, then it has 35 countries with scores less than it.

31 countries divided by 100 gives the particular country a percentile ranking of 31/100 * 100 = 31%.

35 countries divided by 100 gives the particular country a percentile ranking of 35/100 * 100 = 35%.

the percentile ranking went from 31 to 35 which would be an increase of (35% - 31%) / 31% = 4% / 31% * 100 = a percent increase of 12.9% approximately.

i looked on the HDI website and couldn't find anything that showed improvement in position.

they do rank the countries by their index, so if the country went from position 69 to 65, theh it's HDI must have increased by a certain amount.

the improvement could then be calculated off the improvement in the HDI itself, and not the improvement in the position

here's some date from the HDI website

http://hdr.undp.org/en/composite/trends

here's a reference on how to calculate percentile ranking.

http://www.statisticshowto.com/probability-and-statistics/percentiles-rank-range/

here's the main HDI website.

http://hdr.undp.org/en/content/human-development-index-hdi

to look for improvement in the HDI of a country, i would look to the HDI itself.

what was the HDI when the country was in the 69th position?

what was the HDI when the country was in the 65% position?

calculate the percent increase in the HDI from that?

if you just want to work off the position itself, you will need to understand the percentile ranking rather than the position.

if the country was in the 35th percentile when it's rank was 69 and is now in the 40th percentile when it's rank is 65, then the percent increase in the percentile ranking is 40/35 = 1.14 rounded to 2 decimal places minus 1 = .14 * 100 = 14% increase in percentile ranking.l

to understand exactly how this might be cone, you would need to contact the people at HDI to understand how they would calculate it.

if i had to calculate improvement from the position itself, i would first have to find the number of countries in the index when the particular country was in 69th position and when the country was in 65th position.

the number of countries may be the same but it could also be different.

for example, say 200 countries when the country was in position 69 and 200 countries when the country was in position 65.

the country went from 69 to 65 which is an improvement of 4.

200 - 69 = 131 countries below it out of 200 which gives it a percentile rank of 131/200 = .655 * 100 = 65.5%.

200 - 65 = 135 COUNTRIES BELOW IT OUT OF 200 which gives it a percentile rank of .675 * 100 = 67.5%.

67.5% / 65.5% = 1.03 - 1 * 100 = 3% improvement in percentile rank.

now suppose 200 countries when the country was in position 69 and only 100 countries when the country was in position 65.

the calculation would be different.

you would get a percentile rank of 200 - 69 = 131 / 200 = 65.5% when the country was in positon 69.

you would get a percentile rank of 100 - 65 = 35 / 100 = 35% when the country was in position 65.

the percentile rank would then have decreased by a large amount based on the fact that the total number of countries in the index dropped dramatically.

this is not realistic, but goes to show you can't just go on position change alone.

i would not expect the number of countries in the index to change dramatically and will probably remain the same or change by a very small amount as new countries get added to the index and old countries get removed from the index for whatever reason.

relative increase compared to other countries should probably be measured by percentile changes.

absolute increase within the country itself should probably be measured by change in the HDI rating itself.

for more information, consulte the people who do the calculations that provide the overall HDI rank.